# LAVIB: A Large-scale Video Interpolation Benchmark

### v1

@@ Abstract
This paper introduces a LArge-scale Video Interpolation Benchmark (LAVIB) for the low-level video task of video frame interpolation (VFI). LAVIB comprises a large collection of high-resolution videos sourced from the web through an automated pipeline with minimal requirements for human verification. Metrics are computed for each video’s motion magnitudes, luminance conditions, frame sharpness, and contrast. The collection of videos and the creation of quantitative challenges based on these metrics are under-explored by current low-level video task datasets. In total, LAVIB includes 283K clips from 17K ultra-HD videos, covering 77.6 hours. Benchmark train, val, and test sets maintain similar video metric distributions. Further splits are also created for out-of-distribution (OOD) challenges, with train and test splits including videos of dissimilar attributes. 111Our dataset is available at https://alexandrosstergiou.github.io/datasets/LAVIB.
@@ 1 Introduction
Long uncompressed video streams capture events over varying motion intensities, light conditions, and color dynamic ranges. Although loading and storing individual videos is rudimentary, processing and reading large volumes of videos can bottleneck their availability. The high-volume transfer of videos with large filesizes can also result in bandwidth overheads and long decoding times. Low-level vision tasks such as Video Frame Interpolation (VFI) [2, 9, 15, 18, 20, 28, 38, 39, 41, 46, 67], Video Super-Resolution (VSR) [4, 12, 16, 17, 21, 26, 29, 50, 58, 60], and Video Denoising (VD) [13, 30, 31, 49, 54, 57, 59] aim to address such challenges by enabling the storage and stream of lower-resolution, lower-frame-rate, compressed videos. Despite the wide application of such approaches to adjacent tasks such as localization and mapping [24, 63], object tracking [66], novel view synthesis [40, 48], and slow-motion video generation [18, 19], existing datasets for low-level video tasks [1, 37, 38, 43, 51, 52, 55, 56, 57, 64] contain videos of low total durations, with a small number of frames per video. With the exception of [64] most of these datasets only include either a few hundreds [1, 38, 43, 57] or thousands [37, 51, 52, 55, 56] of videos with limited variations in the motions, luminance, and object-level sharpness. To address this gap, this paper introduces a LArge Video Interpolation Benchmark (LAVIB), for learning to extend high-resolution videos across varying motion, blur, luminance, and contrast settings. LAVIB is built on per-frame metrics that quantitatively measure motion magnitudes, frame sharpness, video contrast, and overall luminance. In Fig. 1, LAVIB videos are visualized over axes corresponding to the metrics used.
@@ 2 Related works
LAVIB is compared in Table 1 to adjacent video datasets over different statistics. Dataset statistics include the number of videos and total running times. Video statistics relate to video information such as the resolution and frame rate. Average video metrics provide metrics on the variance of motions, lighting conditions, and frame sharpness. Definitions of the metrics are detailed in §3. LAVIB has threefold more videos than [64] and equally larger total video running time than [55]. The difference in LAVIB video conditions and recording sensors is reflected by the high variance across metrics in Table 1. With the exception of [51], tailored for videos of fast motions with high optical flow magnitude, LAVIB has the highest variance per metric across datasets.
| Dataset | Dataset statistics | Video statistics | Video metrics | |||||
@@ 3 LAVIB statistics
Four statistics are used to obtain segments, create splits, and define challenges. An overview is shown in Table 2 with the number of videos with the highest/lowest metrics reported.
Frame-pair motion. A significant challenge for VFI methods is learning to model cross-frame motion consistency in videos. Thus, the proposed dataset includes videos of diverse magnitudes; both high camera or object motion, and more static scenes. Motion magnitudes can be quantified with dense optical flow. FlowFormer [14] is used on each frame pair resulting in 598 frame pairs per video. The spatial resolution of videos is reduced by to fit frames in memory. The Averaged Flow Magnitude (AFM) is defined by spatio-temporally averaging optical flow. AFM variances are reported for all datasets in Table 1.
@@
As shown in Table 2, videos selected for all splits are balanced across metrics. This is done explicitly for the main benchmark and not the OOD challenges.
@@ 4 LAVIB pipeline
The video selection pipeline includes several stages for the collection, extraction, and set assignment. Initially, videos are searched on YouTube by textual prompts designed to return relevant videos with 4K resolutions and 60 frames per second as overviewed in §4.1. Sourced videos are cropped to 10-second clips standardizing their durations and improving processing speeds in further steps of the collection pipeline. Segments with high motion magnitudes are selected from the clips and are spatially cropped based on the AVL, ARL, and ARMS statistics detailed in §4.2. Dataset splits are balanced between the four statistics, with OOD splits created by assigning videos with the highest average metric at the test or train set. The dataset pipeline is flexible and can be scaled over large numbers of videos, requiring manual input only at a few points.
4.1 Web-Crawling and video selection
@@
Videos should be in the wild. Retrieved videos should vary by lighting conditions, motions, and scenes. They should also be recorded with different sensors. Sensor types depend strongly on video themes; e.g. action cameras are more common for capturing fast-paced scenes in contrast to DSLR cameras. Conditions are added in the format; ‘rainy walk in New York’ or ‘night drive’. Some text prompts are also designed to include specific equipment such as ‘GoPro Hero10’ or ‘iPhone 13 Pro’.
@@
4.2 Segment selection and assignment
In total, 667 hours of footage are collected over the project’s 31-month duration. This initial list contained videos of hour-long to minute-long durations. To standardize their durations, 10-second clips are sampled manually over different interval steps. Clips are extracted consecutively for videos less than 5 minutes. For the rest of the videos; 10-second sampling intervals are used for videos with durations between 5-30 minutes, 2-minute intervals for videos between 30 minutes to an hour, and 10-minute intervals for videos longer than an hour. This selection resulted in a total of 34,408 clips. Clips from the same video are bound to include similarities. To account for this and inspired by [68], similarities between clips from the same video are measured metrically by their embedding space distance with highly similar clips being dropped. MViTv2-B [10] is used as the clip encoder from which a similarity matrix between final layer embeddings is created by computing the L2 distance over clip pairs. Clips with an average (row-wise) L2 distance below the entire matrix’s average are dropped. This step resulted in the selection of 17,204 clips.
@@ 4.2 Segment selection and assignment
Temporal Segment Selection. Segments are compared and in turn selected by their average motion magnitudes. Primarily static segments are dropped as they are less relevant to VFI tasks with minimal pixel and object tracking requirements. FlowFormer [14] calculates flow maps over pairs of frames with magnitudes averaged spatially. Each 10-second sequence is temporally augmented to obtain all available 1-second clips. Clips with the highest 20% magnitudes are selected. This strategy was chosen as it worked well in a small-scale setting when manually examining a set of 1,000 clips.
Spatial Segment Selection. The selected high-resolution clips cannot directly fit into the memory of most current GPUs. Thus, as commonly addressed in the literature [35, 38, 51, 64] the number of videos is curated with the additional selection of tubelets. Each clip is divided into four tubelets for the top right, top left, bottom right, and bottom left parts of the videos. Tubelets are selected by the ALV, ARL, ARMS, and each tubelet’s AFM. Starting from low/high values for each metric, 80% of the tubelets are retained. This avoids oversampling from values close to the mean of metrics. Instead, tubelets with more challenging settings are selected.
Assignment to splits. All train/val/test splits are constructed with a 65-15-20% split. DUPLEX [53] selection is used for balancing split statistics. Videos with the largest pair-wise distance by their metrics are initially selected. In turn, videos are iteratively assigned to sets given their distance from the previously selected videos. A detailed overview of the algorithm is provided in §1. Recall that the OOD sets need to be imbalanced across statistics so this is specific to the benchmark splits.
@@ 5 Benchmarks
Baselines. LAVIB contains 188,644 1-second videos for training, 41,345 videos for validation, and 53,494 videos for testing. Benchmark results are reported in §5.1 across settings. For the baselines, triplets of frames are defined similarly to [6, 55, 64] for single-frame interpolation with a total of 5.7M triplets. In the multi-frame interpolation settings in §5.2, septuplets are also used resulting in a total of M groups. In §5.3, the video metrics are used to create unbalanced dataset splits. For each of the four metrics, two challenges are created by sampling videos with either high/low values and assigning them to the train/test sets. Qualitative results for all three models are shown in §5.4.
@@ 5.1 Baseline results
Baselines. Table 3 reports SSIM, PSNR, and LPIPS scores on both LAVIB val and test sets across three training settings; pre-training on Vimeo-90K, fine-tuning on a joint set from Xiph [35, 38] and X4K1KFPS [51] of exclusively 4K videos, and pre-training with LAVIB. FLAVR [22] is used as the baseline model due to its fast processing times, strong results, and open-source codebase. Finetuning on Xiph + X4K1KFPS suffers as both datasets are small in size although they are sourced by videos with the same resolution as LAVIB. Pre-training only on Vimeo-90K slightly improves results. Pre-training on LAVIB gives the best performance overall increasing PSNR, and SSIM by +1.08 and +0.015 on average on both sets.
Generalization to related small-scale datasets. VFI benchmarks include multiple datasets [35, 38, 51, 64]. LAVIB is unique in having the largest number of diverse videos of both high resolution and high frame rates. The generalization benefits of using LAVIB as the pre-training dataset are compared to the previously widely-used Vimeo-90K dataset [64]. Table 6 shows performance improvements in the test set of Vimeo-90K when the model is trained on LAVIB. Similar score increases are also observed on the Xiph4K and X4K1KFPS test sets with +1.23 and +1.19 improvements on the PSNR. LAVIB’s large variance across videos enables learning VFI over different conditions which can benefit performance in smaller task-specific benchmark datasets.
Multi-metric results. As human judgment of the perceptual quality depends on high-order image structures and context [33, 61] an ensemble of metrics is reported in Table 7 to provide a complete evaluation of methods’ performance on the LAVIB test set. In addition to standard quality metrics, scores over recently-proposed metrics including DISTS [8], Watson-DFT [7], VSFA [27], and VFIPS [11] are also reported. Across statistics, both EMA-VFI and FLAVR perform comparably. A decrease in performance is observed with RIFE as its limited complexity can not adequately address VFI with large variations in settings across videos. Compared to FLAVR, the PSNR and SSIM scores decrease by -5.56 and -1.10 respectively, and the LPIPS loss increases from 0.029 to 0.146.
@@ 5.2 Multi-frame interpolation results
This section ablates the number of frames interpolated and evaluated over different schemes with; interpolation being equivalent to interpolating 30fps videos to 60fps, interpolating 20fps to 60fps, and interpolating 15fps to 60fps. Triplets and septuplets of frames as input are also ablated. Results are reported in Table 8. FLAVR trained on Vimeo-90K is used as a baseline in all settings.
@@ 5.3 OOD Challenges
Low/High AFM. As shown in Table 9(a), VFI models cannot effectively interpolate frames when trained on videos with low motion magnitudes. Compared to the benchmark results in Table 7 a -2.64 and -0.04 drop is observed for PSNR and SSIM. The embedding distance to ground truth frames also increases by +9.825e. In contrast, when models are trained on high motion magnitudes, VFI is easier for the target domain of primarily low magnitudes. The imbalance in performance shows the sensitivity of current models to the motion magnitudes of the training data.
Low/High ALV. Sharpness-based comparisons are reported in Table 9(b). Testing on low-sharpness settings is more challenging for VFI models as object edges are more difficult to define. However, models trained on low-sharpness videos can interpolate high-sharpness videos with an average +2.93 and +0.023 increase in the PSNR and SSIM scores compared to the low-to-high task.
Low/High ARMS. Results on contrast-based OOD challenges are presented in Table 9(c). The domain gap between these two settings is significant. Training on low contrast shows robustness when the domain shifts to high contrast at testing. However, the same generalization is not observed when models are trained on high-contrast videos and tested on low-contrast VFI. Compared to low-contrast training and high-contrast testing, high-to-low shows a -1.65 drop in PSNR.
Low/High ARL. Table 9(d) reports performances over brightness settings. Overall, models from either setting show comparable performance and generalization robustness to the target domain. Minor performance improvements are shown for the high to low task with high luminance training being more effective in cross-domain generalization.

### v2

@@ Abstract
This paper introduces a LArge-scale Video Interpolation Benchmark (LAVIB) for the low-level video task of Video Frame Interpolation (VFI). LAVIB comprises a large collection of high-resolution videos sourced from the web through an automated pipeline with minimal requirements for human verification. Metrics are computed for each video’s motion magnitudes, luminance conditions, frame sharpness, and contrast. The collection of videos and the creation of quantitative challenges based on these metrics are under-explored by current low-level video task datasets. In total, LAVIB includes 283K clips from 17K ultra-HD videos, covering 77.6 hours. Benchmark train, val, and test sets maintain similar video metric distributions. Further splits are also created for out-of-distribution (OOD) challenges, with train and test splits including videos of dissimilar attributes. 111LAVIB is accessible at https://alexandrosstergiou.github.io/datasets/LAVIB.
@@ 1 Introduction
Long uncompressed video streams capture events over varying motion intensities, light conditions, and color dynamic ranges. Although loading and storing individual videos is rudimentary, processing and reading large volumes can bottleneck availability. The high-volume transfer of videos with large filesizes can also result in bandwidth overheads and long decoding times. Low-level vision tasks such as Video Frame Interpolation (VFI) [3, 10, 16, 19, 21, 30, 41, 42, 44, 50, 75], Video Super-Resolution (VSR) [5, 13, 17, 18, 22, 28, 31, 54, 63, 65], and Video Denoising (VD) [14, 32, 33, 53, 58, 61, 64] aim to address such challenges by enabling the storage and stream of lower-resolution, lower-frame-rate, compressed videos. Despite the wide application of such approaches to adjacent tasks such as localization and mapping [26, 71], object tracking [74], novel view synthesis [43, 52], and slow-motion video generation [19, 20], existing datasets for low-level video tasks [2, 40, 41, 46, 55, 56, 59, 60, 61, 72] contain short videos, with a small number of frames per video. With the exception of [72] most of these datasets only include either a few hundreds [2, 41, 46, 61] or thousands [40, 55, 56, 59, 60] of videos with limited variations in the motions, luminance, and object-level sharpness. To address this gap, this paper introduces a LArge Video Interpolation Benchmark (LAVIB), for learning to interpolate high-resolution videos across varying motion, blur, luminance, and contrast settings. LAVIB is built on per-frame metrics that quantitatively measure motion magnitudes, frame sharpness, video contrast, and overall luminance. In Fig. 1, LAVIB videos are visualized over axes corresponding to the metrics used.
@@ 2 Related works
LAVIB is compared in Tab. 1 to adjacent video datasets over different statistics. Dataset statistics include the number of videos and total running times. Video statistics relate to video information such as the resolution and frame rate. Average video metrics provide metrics on the variance of motions, lighting conditions, and frame sharpness. Definitions of the metrics are detailed in §3. LAVIB has threefold more videos than [72] and equally larger total video running time than [59]. The difference in LAVIB video conditions and recording sensors is reflected by the high variance across metrics in Tab. 1. With the exception of [55], tailored for videos of fast motions with high optical flow magnitude, LAVIB has the highest variance per metric across datasets.
| Dataset | Dataset statistics | Video statistics | Average video metrics | |||||
@@ 3 LAVIB statistics
Four statistics are used to obtain segments, create splits, and define challenges. An overview is shown in Tab. 2 with the number of videos with the highest/lowest metrics reported.
Frame-pair motion. A significant challenge for VFI methods is learning to model the cross-frame motion consistency of videos. Thus, the proposed dataset includes videos of diverse magnitudes; both high camera or object motion, and more static scenes. Motion magnitudes can be quantified with dense optical flow. FlowFormer [15] is used on each frame pair resulting in 598 frame pairs per video. The spatial resolution of videos is reduced by to fit frames in memory. The Averaged Flow Magnitude (AFM) is defined by spatio-temporally averaging optical flow. AFM variances are reported for all datasets in Tab. 1.
@@
As shown in Tab. 2, videos selected for all splits are balanced across metrics. This is done explicitly for the main benchmark and not the OOD challenges.
@@ 4 LAVIB pipeline
The video selection pipeline includes several stages for the collection, extraction, and set assignment. Initially, videos are searched on YouTube by textual prompts designed to return relevant videos with 4K resolutions and 60 frames per second as overviewed in §4.1. Sourced videos are cropped to 10-second clips standardizing their durations and improving processing speeds in further steps of the collection pipeline. Segments with high motion magnitudes are selected from the clips and are cropped to tubelets by their AVL, ARL, ARMS, and AFM statistics as detailed in §4.2. Dataset splits are balanced between the four statistics, with OOD splits created by assigning videos with the highest average metric at the test or train set. The dataset pipeline is flexible and can be scaled over large numbers of videos, requiring manual input only at a few points.
4.1 Video web-crawling
@@
Videos should be in the wild. Retrieved videos should vary by lighting conditions, motions, and scenes. They should also be recorded with different sensors. Sensor types depend strongly on video themes; e.g. action cameras are more common for capturing fast-paced scenes in contrast to DSLR cameras. Conditions are added in the format; rainy walk in New York or night drive. Some text prompts are also designed to include specific equipment such as GoPro Hero10 or iPhone 13 Pro.
@@
4.2 Segment selection and split assignment
In total, 667 hours of footage are collected over the project’s 31-month duration. This initial list contained videos of hour-long to minute-long durations. To standardize their durations, 10-second clips are sampled manually over different interval steps. Clips are extracted consecutively for videos less than 5 minutes. For the rest of the videos; 10-second sampling intervals are used for videos with durations between 5-30 minutes, 2-minute intervals for videos between 30 minutes to an hour, and 10-minute intervals for videos longer than an hour. This selection resulted in a total of 34,408 clips. Clips from the same video are bound to include similarities. To account for this and inspired by [76], similarities between clips from the same video are measured metrically by their embedding space distance with highly similar clips being dropped. MViTv2-B [11] is used to encode clips and to create a similarity matrix based on the L2 distance of the final layer embeddings. Clips with an average (row-wise) L2 distance below the entire matrix’s average distance are dropped. This step resulted in the selection of 17,204 clips.
@@ 4.2 Segment selection and split assignment
Temporal segment selection. Segments are compared and selected by their AFM. This selection aims to drop primarily static segments as they are less relevant to VFI tasks with minimal pixel and object tracking requirements. FlowFormer [15] is used to calculate AFM over pairs of frames by spatially averaging flows. Each 10-second sequence is temporally augmented to obtain all available 1-second clips. Clips with the highest 20% magnitudes are selected. This strategy was chosen as it worked well in a small-scale setting when manually examining a set of 1,000 clips.
Spatial segment selection. The selected high-resolution clips cannot directly fit into the memory of most current GPUs. Thus, as commonly addressed in the literature [38, 41, 55, 72] the number of videos is curated with the additional selection of tubelets. Each clip is divided into four tubelets by a grid. ALV, ARL, ARMS, and AFM are computed for each tubelet. 80% of the tubelets are retained by selecting from the low/high values per metric in succession, leaving out the middle 20%. This avoids oversampling from values close to the mean of metrics. Instead, tubelets with more challenging settings are selected.
[…]
@@ 5 Benchmarks
Baselines. LAVIB contains 188,644 1-second videos for training, 41,345 videos for validation, and 53,494 videos for testing. Benchmark results are reported in §5.1 across settings. For the baselines, triplets of frames are defined similarly to [7, 59, 72] for single-frame interpolation with a total of 5.7M triplets. In the multi-frame interpolation settings in §5.2, septuplets are also used resulting in a total of M groups of frames. Ablations on varying video resolutions are presented in §5.3. In §5.4, the video metrics are used to create unbalanced dataset splits. For each of the four metrics, two challenges are created by sampling videos with either high/low values and assigning them to the train/test sets. Qualitative results for all three models are shown in §5.5.
@@ 5.1 Baseline results
††footnotetext: ∗Inhouse evaluation from author provided model.Baselines. Tab. 3 reports SSIM, PSNR, and LPIPS scores on both LAVIB val and test sets across three training settings; pre-training on Vimeo-90K, fine-tuning on a joint set from Xiph [38, 41] and X4K1KFPS [55] of exclusively 4K videos, and pre-training with LAVIB. FLAVR [23] is used as the baseline model due to its fast processing times, strong results, and open-source codebase. Finetuning on Xiph + X4K1KFPS suffers as both datasets are small in size although they are sourced by videos with the same resolution as LAVIB. Pre-training only on Vimeo-90K slightly improves results. Pre-training on LAVIB gives the best performance overall increasing PSNR, and SSIM by +1.08 and +0.015 on average on both sets.
Generalization to related small-scale datasets. VFI benchmarks include multiple datasets [38, 41, 55, 72]. LAVIB is unique in having the largest number of diverse videos of both high resolution and high frame rates. The generalization benefits of using LAVIB as the pre-training dataset are compared to the previously widely-used Vimeo-90K [72]. Tab. 6 shows performance improvements in the test set of Vimeo-90K when the model is trained on LAVIB. Similar score increases are also observed for the Xiph4K and X4K1KFPS test sets in Tabs. 6 and 6 with +1.23 and +1.19 improvements on the PSNR. LAVIB’s large variance across videos enables learning VFI over different conditions which can benefit performance in smaller domain-specific benchmark datasets.
Multi-metric results. As human judgment of the perceptual quality depends on high-order image structures and context [36, 67], an ensemble of metrics is reported in Tab. 7 to provide a complete evaluation of each methods’ performance on the LAVIB test set. In addition to standard quality metrics, scores over recently-proposed metrics including DISTS [9], Watson-DFT [8], VSFA [29], and VFIPS [12] are also reported. Across statistics, both EMA-VFI and FLAVR perform comparably. A decrease in performance is observed with RIFE as its limited complexity can not adequately address VFI with large variations in settings across videos. Compared to FLAVR, the PSNR and SSIM scores decrease by -5.56 and -1.10 respectively, and the LPIPS loss increases from 0.029 to 0.146.
@@
| Model | ||||
| PSNR | SSIM | PSNR | SSIM | |
| FLAVR | 33.21 | 0.978 | 33.44 | 0.981 |
| Train set | ||||||
| PSNR | SSIM | LLIPS | PSNR | SSIM | LLIPS | |
| Video90K | 30.14 | 0.943 | 4.638e-2 | 32.10 | 0.963 | 3.947e-2 |
| LAVIB | 32.57 | 0.965 | 3.781e-2 | 33.44 | 0.981 | 2.934e-2 |
| Model | |||||||||
| PSNR | SSIM | LLIPS | PSNR | SSIM | LLIPS | PSNR | SSIM | LLIPS | |
| EMA-VFI | 32.26 | 0.954 | 4.130e-2 | 33.01 | 0.972 | 3.211e-2 | 33.14 | 0.978 | 3.105e-2 |
| FLAVR | 32.57 | 0.965 | 3.781 | 33.28 | 0.973 | 3.086 | 33.44 | 0.981 | 2.934e-2 |
@@ 5.2 Multi-frame interpolation results
This section ablates the number of frames interpolated and evaluated over different schemes with; interpolation being equivalent to interpolating 30fps videos to 60fps, interpolating 20fps to 60fps, and interpolating 15fps to 60fps. Triplets and septuplets of frames as input are also ablated. Results are reported in Tab. 8. FLAVR trained on Vimeo-90K is used as a baseline in all settings.
@@
Varying frame sampling. LAVIB’s standardized 60fps also enables works to explore VFI over more challenging settings with multiple temporal resolutions. Tab. 10 reports performance on 30fps targets created by sampling every 2 frames to form triplets. Results show consistency between densely sampling frames sequentially () and sampling with a step of 2 ().
5.3 Varying frame resolution results
An important factor for VFI is the clarity of the objects. Different computational budgets can limit availability in training schemes and memory use.
Resolutions across models. Results on different training set resolutions are reported in Tab. 11. As in [16, 23, 75], is the standard resolution used for training all models. A proportional decrease in performance is observed at lower resolutions. However, these reductions remain small with an average / in PSNR/SSIM when using and / when using . Thus, LAVIB can be a suitable benchmark for evaluating low-compute VFI models.
Frame resolutions across training schemes. Tab. 10 reports performances across varying resolutions with different dataset training sets. Compared to the LAVIB-trained model, performance degrades significantly at lower resolutions with the smaller and less diverse Vimeo-90K. The large and varying LAVIB training set can be an effective alternative for training on lower compute resources in which full-resolution videos do not fit in memory.
@@ 5.4 OOD Challenges
Low/High AFM. As shown in Tab. 12(a), existing VFI models cannot effectively interpolate frames when trained on videos with low motion magnitudes. Compared to the benchmark results in Tab. 7 a and drop is observed for PSNR and SSIM. The embedding distance to ground truth frames also increases by +9.825e. In contrast, when models are trained on high motion magnitudes, VFI is easier for the target domain of primarily low magnitudes. The imbalance in performance shows the sensitivity of current models to the motion magnitudes of the training data.
Low/High ALV. Sharpness-based comparisons are reported in Tab. 12(b). Testing on low-sharpness settings is more challenging for VFI models as object edges are more difficult to define. However, models trained on low-sharpness videos can interpolate high-sharpness videos with an average +2.93 and +0.023 increase in the PSNR and SSIM scores compared to the low-to-high task.
Low/High ARMS. Results on contrast-based OOD challenges are presented in Tab. 12(c). The domain gap between these two settings is significant. Training on low contrast shows robustness when the domain shifts to high contrast at testing. However, the same generalization is not observed for the inverse with models trained on high-contrast videos and tested on low-contrast VFI. Compared to low-to-high ARMS, high-to-low ARMS shows a -1.65 drop in PSNR.
Low/High ARL. Tab. 12(d) reports performances over brightness settings. Overall, models from either setting show comparable performance and generalization robustness to the target domain. Minor performance improvements are shown for the high to low task with high luminance training being more effective in cross-domain generalization.